In a test of the effectiveness of garlic for lowering​ cholesterol,48 subjects were treated with garlic in a processed tablet form. Cholesterol levels were measured before and after the treatment. The changes​(beforeminus−​after)in their levels of LDL cholesterol​ (in mg/dL) have a mean of 5.4and a standard deviation of 17.7.Construct a 90​%confidence interval estimate of the mean net change in LDL cholesterol after the garlic treatment. What does the confidence interval suggest about the effectiveness of garlic in reducing LDL​ cholesterol?

What is the confidence interval estimate of the population mean

muμ​?
Answer #1

Degrees of freedom = n - 1 = 48 - 1 = 47

For 47 degrees of freedom and 90% confidence interval,

Critical t value = 1.678

Hence,

90% confidence interval will be:

5.4 1.678 * 17.7/V48

5.4 \pm 4.29

(1.11, 9.69)

Since the confidence interval contains only positive values, it suggests that garlic is effective in reducing :D: cholesterol.